shanti sweets stall was placing an order for making cardboard boxes for packing their sweets.Two sizes of boxes were required.The bigger of dimensions 25cm*20cm*5cm and the smaller of dimensions 15cm*12cm*5cm.For all the overlaps, 5% of the total surface area is required extrs.If the cost of the cardboard is rs 4 for 1000 cm2 , find the cost of cardboard required for supplying 250 boxes of each kind.
Answer by mananth(16946)

25cm*20cm*5cm -------------- 250 boxes
surface areaof 1 box = 2(25*20)+2(20*5)+2(5*25)=1450 sq.ft
for 250 boxes=362500 cm^2
5% extra
=362500*1.05=380625 cm^2
-----------
15cm*12cm*5cm.----------------- 250 boxes
surface area = 2(15*12)+2(12*5)+2(5*15)=630 cm^2
250 boxes = 250*630=157500 cm^2
5% extra =1.05*157500=165375 cm^2
Total card board required=362500+165375=527875 cm^2
for 1000 cm^2 Rs.4
for 527875 cm^2= ?
Rs. 2111.50
m.ananth@hotmail.ca